top of page

2018 data update (28/n)

[Orig: June 18, 2018]

Hi everyone,

My apologies for the extended silence over here. LCO had changed the instrument configuration in the beginning of June. This lead to a several issues in the standard data pipelines, and it took us a while to work through things. Data taken during the last week or so is updated on this new graph:

The gray vertical line indicates the time of the change in configuration. There are still some remaining issues that we need to address for data taken primarily during the first week of June (where you see a gap in data) and hopefully we will be able to update this soon with those measurements. Also, all data taken after this time have an unknown vertical offset. As such, these measurements are set to be consistent with the flux at the end of May until we can better assess what, if anything, had changed during the gap.

Bottom line: the star's brightness is still slightly above normal levels ("normal" is defined to be the dotted horizontal line, set to unity at the time just before Elsie in 2017), and no large dips have been observed since 2018 March.

More later!

~Tabby and team

PS: These observations are happening because of the wonderful backers of our 2016 Kickstarter project. The Kickstarter campaign has ended, but we are still accepting donations to purchase additional observing time on the LCO 0.4m network. Thanks in advance for your support!

Featured Posts
Recent Posts
Search By Tags
No tags yet.
Follow Us
  • Facebook Basic Square
  • Twitter Basic Square
  • Google+ Basic Square
bottom of page